# The Distribution of Atomic Hydrogen in the Host Galaxies of FRBs

### v2

@@ Abstract
We probe the atomic hydrogen (Hi) emission from the host galaxies of fast radio bursts (FRBs) to investigate the emerging trend of disturbance and asymmetry in the population. Quadrupling the sample size, we detect 13 out of 14 new hosts in Hi, with the single non-detection arising in a galaxy known to be transitioning towards quiescence. With respect to typical local Universe galaxies, FRB hosts are generally massive in Hi (HI ), which aligns with previous studies showing that FRB hosts also tend to have high stellar masses and are star-forming. However, they span a broad range of other Hi derived properties. In our independent sample of repeater hosts, we observe a statistically insignificant preference towards lower Hi masses compared to non-repeater hosts, similar to the low-significance trend toward lower stellar masses previously reported. Using visual inspection alongside various asymmetry metrics, we identify four unambiguously settled host galaxies, demonstrating for the first time that a disturbed Hi morphology is not a universal feature of FRB host galaxies. However, we find another six that show clear signs of disturbance, and three which require deeper or more targeted observations to reach a conclusion; this brings the confirmed ratio of disturbed-to-settled FRB hosts to 11:4. Given that roughly a 1:1 ratio is expected for random background galaxies of similar type, our observed ratio yields a p-value of 0.065. Unlike earlier indications based on smaller samples, this no longer crosses the conventional threshold for statistical significance, though is still near enough to hint at a legitimate excess of disturbance among FRB hosts. Thus, an even larger sample size of FRB hosts observed in Hi is required to fully clarify whether the trend is genuine or still a consequence of low-number statistics — a sample that upcoming data releases are well positioned to provide.
@@ 1 Introduction
Fast radio bursts (FRBs) are energetic pulses of radio emission that generally last for a few milliseconds in timescale (Lorimer2007). Whilst their extragalactic origins are well established (Chatterjee2017; Bannister2019), their generation mechanisms remain unclear, despite significant advances in the field. Dozens of progenitor models have been proposed to explain the phenomena, from outbursting neutron stars to interacting black holes (for reviews, see Platts2019; Petroff2019). The former, specifically young, highly magnetised neutron stars or magnetars, are favoured within the community due to observations of energetic FRBs found to be embedded within persistent radio sources (PRSs) (e.g. Metzger2017). This link was strengthened in 2020 with the detection of an FRB from a magnetar within our galaxy (Andersen2020; Bochenek2020), although its isotropic-equivalent energy falls nearly 30 times lower than the next weakest FRB.
@@
Since their discovery, the population of detected FRBs has rapidly expanded, primarily due to dedicated observations from the Australian Square Kilometre Array Pathfinder (ASKAP; ASKAP; Bannister2019; Shannon2024), the Canadian Hydrogen Intensity Mapping Experiment (CHIME; CHIME; ChimeC1), MeerKAT (MeerKAT; MeerTRAP), and the Deep Synoptic Array (DSA; Kocz2019; DSA; Connor2025). Crucially, improvements to spatial resolution have allowed many FRBs to be localised to arcsecond or better accuracy, allowing for deep follow up photometry and spectroscopy of their host galaxies and environments (e.g. Bannister2019; Heintz2020; Mannings2021; Bhandari2022; Gordon2023; Ryder2023; Sharma2024; Amiri2025). Such localisations are key to revealing the environments of FRBs and unlocking their potential for cosmological studies (e.g. McQuinn2014; Macquart2020; James2022; Connor2025).
@@
As of April 2025, over 100 host galaxies have been identified, spanning redshifts out to 1.33. These galaxies tend to be star-forming spiral galaxies, though most other galactic parameters such as stellar mass and mass-weighted age are highly varied (Bhandari2022; Gordon2023). These observations broadly support a magnetar progenitor hypothesis, as magnetars are expected to trace star formation through a core-collapse supernovae generation channel (Metzger2017). However, a number of counterexamples, such as a localisation to a globular cluster (Kirsten2022), and several to quiescent galaxies (Eftekhari2025; Shah2025), muddy the waters. Furthermore, no statistically significant trend differing the hosts of repeating and apparently non-repeating FRBs has been observed (Gordon2023; Sharma2024), despite significant evidence suggesting fundamental differences in their respective origins (ChimeC1).
@@
Beyond analysis of the optical output of these host galaxies, a number of studies have been conducted on their atomic hydrogen (Hi) gas emission (Michalowski21; Kaur2022; Glowacki_wallaby; Lee-waddell23; Glowacki_hi_frb_whats_your_z). At redshifts greater than 0.1, the Hi line becomes weak and contaminated by RFI; with much of the localised FRB population lying beyond this redshift, only six host galaxies (excluding the Milky Way) have thus far been probed. All six of these galaxies have been detected in Hi, an expected result considering the star-forming nature of most hosts as derived from optical observations. Interestingly, a trend has emerged, with five of six of these detections exhibiting either a strongly asymmetric Hi spectrum or a highly disturbed spatial gas distribution. The single apparently symmetric profile, that of FRB 20211127I, was analysed with a low resolution and low signal-to-noise ratio (SNR) ASKAP observation, and thus its lack of disturbance was considered inconclusive (Glowacki_wallaby).
@@
Such features, which are well documented in the literature (e.g. Holwerda2011; Espada2011; Bok2019; Reynolds2020), are attributed to recent or ongoing disruptions to the galactic environment. The gaseous disk is sensitive to both external and internal perturbative mechanisms, including tidal interactions and mergers, ram pressure stripping, and feedback from supernovae or active galactic nuclei (AGN; Yu2022). Disturbed Hi distributions are not uncommon; many studies have found that a significant fraction of galaxies exhibit some sign of disturbance. Historical studies using single dish instruments focused on the asymmetry of the 1D profile; early estimates found that at least 50% of galaxies exhibit spectra with some level of asymmetry (Richter94; Haynes1998; Matthews1998). However, more recent studies use more stringent cutoffs; including these results (Espada2011; Watts2020; Reynolds2020; Yu2022) and reevaluating the historical ones find anywhere between 10-40% of galaxies exhibit asymmetric Hi line profiles.
@@
In comparison, few studies have investigated similar properties in the spatial distribution of Hi resolved galaxies (Holwerda2011; vanEymeren). Reynolds2020LVHIS recently probed 120 galaxies from the LVHIS (LVHIS), VIVA (VIVA), and HALOGAS (HALOGAS) surveys, finding similar rates of asymmetry. However, with a number of more advanced interferometric observatories currently in the process of conducting large Hi surveys of resolved galaxies — such as the Widefield ASKAP L-band Legacy All-sky Blind surveY (WALLABY, WALLABY) on ASKAP and the Hi project within the MeerKAT International GigaHertz Tiered Extragalactic Exploration survey (MIGHTEE-HI, MIGHTEEHI) — many more observations will be made to properly pin down morphological asymmetry in Hi disks.
@@
Overall, the exact fraction of disturbed galaxies measured by each study is highly dependent on the samples used (e.g., isolated vs galaxy pairs), the method of quantification, and the point of cutoff in these methods above which a profile is considered disrupted. However, it can generally be said that disturbances are common, and are likely present in up to half of the galaxy population.
@@
Thus, given the low number of hosts probed, the early trend towards asymmetry in the population — with a ratio of 5:0 or 5:1 — is not necessarily significant at this stage. However, if this trend perseveres as more host galaxies are observed in Hi, it may indicate a connection between FRBs and a recent enhancement of star formation due to galactic interaction (Michalowski21). Indeed, studies over the past few decades have identified higher star formation rates in merging galaxies (e.g. Alonso2004; Darg2010). Such a scenario would lend further credence to so called ‘fast track’ FRB progenitors such as magnetars, which are expected to trace star formation.
@@ 2.1 FRB Host Galaxy Targets
The FRBs selected for this study are taken from published datasets from the CRAFT survey (Commensal Real-time ASKAP Fast Transients; CRAFT; Bannister2019; Shannon2024), the CHIME/FRB Collaboration (ChimeC1; Amiri2025), and the MeerTRAP survey (More TRAnsients and Pulsars; MeerTRAP; Rajwade2022; Driessen2024).
We place two criteria on targets for inclusion in this study. The first is that the host association is robust, which is evaluated through the PATH framework (Probabilistic Association of Transients to their Hosts; Aggarwal2021). PATH employs a Bayesian approach to calculate the probability a given transient event is associated with a galaxy, based on the event’s localisation and the galaxy’s magnitude and size. Following the convention used in Bhandari2022; Gordon2023, we consider PATH posterior probabilities P(O—) 0.9 to be robust host associations.
Our second criterion is that the redshifts of the target host galaxies are . This is due to both the sensitivity of the Hi line, and also the prevalence of contaminating radio frequency interference (RFI) below 1300 MHz (0.092), largely due to the Global Navigation Satellite System (GNSS).
Based on these cuts, our sample consists of nine CHIME FRBs, three CRAFT FRBs, and two MeerTRAP FRBs. All but one CHIME target (FRB 20200223B) are apparently non-repeating bursts (“non-repeaters” herein), and one of the CRAFT targets (FRB 20211127I) is the FRB previously analysed in Glowacki_wallaby. The details of their host galaxies are presented in Table 1.
@@
FRB 20181220A | CHIME | 2MFGC 17440 | 348.6982 | 48.3421 | 0.02746 | GMRT/FAST | chimefrbs |
FRB 20181223C | CHIME | SDSS J120340.98+273251.4 | 180.9207 | 27.5476 | 0.03024 | GMRT | chimefrbs |
FRB 20190418A | CHIME | SDSS J042314.96+160425.6 | 65.8123 | 16.0738 | 0.07132 | GMRT | chimefrbs |
FRB 20190425A | CHIME | UGC 10667 | 255.6625 | 21.5767 | 0.03122 | GMRT | chimefrbs |
FRB 20200223B† | CHIME | SDSS J003304.68+284952.6 | 8.2695 | 28.8313 | 0.06024 | GMRT/FAST | Ibik2023 |
FRB 20200723B | CHIME | NGC 4602 | 190.1538 | -5.1328 | 0.0085 | MeerKAT | frb20200723b |
FRB 20201123A | MeerTRAP | J173438.35-504550.4 | 263.6596 | -50.7641 | 0.0507 | MeerKAT | Rajwade2022 |
FRB 20210405I | MeerTRAP | 2MASS J1701249-4932475 | 255.3537 | -49.5465 | 0.0656 | MeerKAT | frb20210405 |
FRB 20211127I | CRAFT | WALLABY J131913–185018 | 199.8082 | -18.8378 | 0.04695 | MeerKAT | Glowacki_wallaby |
FRB 20211212A | CRAFT | SDSS J102924.22+012139.2 | 157.3509 | 1.3608 | 0.0715 | MeerKAT | Shannon2024 |
FRB 20231229A | CHIME | UGC 1234 | 26.4658 | 35.1110 | 0.0190 | Arecibo | Amiri2025 |
FRB 20231230A | CHIME | J045109.39+022205.02 | 72.7887 | 2.3681 | 0.0298 | MeerKAT | Amiri2025 |
FRB 20240201A | CRAFT | WISEA J095937.44+140519.4 | 149.9060 | 14.0887 | 0.04273 | MeerKAT/FAST | Shannon2024 |
FRB 20250316A | CHIME | NGC 4141 | 182.4472 | 58.8492 | 0.00635 | Effelsberg | Amiri2025 |
@@ 2.2 Hi Observations & Processing
We first searched for archival Hi datasets, and found observations of the hosts of FRB 20231229A and FRB 20250316A (Amiri2025). Both galaxies have been observed in Hi twice; we take the most recent spectra available in each case, corresponding to an ALFALFA (Arecibo Legacy Fast ALFA; Haynes2018) survey observation and an Effelsberg 100-m observation (Springob2005) respectively. Both of these are at spectral resolutions of 11 km s-1, with single-dish beams encompassing several times the angular size of the galaxy.
A search of the MeerKAT archive hosted by the South African Radio Astronomical Observatory (SARAO) revealed that the coordinates of three targets were contained in archival band observations. The host of FRB 20200723B (frb20200723b) was the target of a 0.5 hour 4K channelisation (corresponding to 44 km s-1resolution) pointing from ”Part 2: 1.28 GHz MeerKAT survey of 12MGS Galaxies in the Southern Hemisphere”, project ID SCI-20220822-LM-01, capture block ID 1671245181. The host of FRB 20201123A (Rajwade2022) was contained in a 1 hour 4K channelisation pointing by the ”Legacy Survey of the Galactic Plane”, project ID SSV-20180505-FC-01, capture block ID 1610418665, a distance of 0.4 from the phase centre. The host of FRB 20231230A (Amiri2025) was contained in a 7.5 hour 4K channelisation pointing from the proposal ”Observation of the galaxy cluster A520”, project ID SCI-20210212-SG-01, capture block ID 1638727749, a distance of 0.9 from the phase centre.
The remaining targets had no other archival data available, and thus we pursued targeted observation with the upgraded Giant Metrewave Radio Telescope (uGMRT; GMRT; uGMRT) and the Five-hundred-meter Aperture Spherical Telescope (FAST; FAST) for northern hemisphere targets, and MeerKAT for southern hemisphere targets. We present summaries of all Hi observations in Table 2.
@@
FRB 20250316A | Effelsbergd | - | 11.965 | - | - | - |
@@
Haynes2018 - observation details are not available.
@@
Springob2005 - observation details are not available.

### v3

@@ Abstract
We probe the atomic hydrogen (Hi) emission from the host galaxies of fast radio bursts (FRBs) to investigate the emerging trend of disturbance and asymmetry in the population. Quadrupling the sample size, we detect 16 out of 17 new hosts in Hi, with the single non-detection arising in a galaxy known to be transitioning towards quiescence. With respect to typical local Universe galaxies, FRB hosts are generally massive in Hi (HI ), which aligns with previous studies reporting that FRB hosts also tend to have high stellar masses and are star-forming. However, they span a broad range of other Hi derived properties. Using visual inspection alongside various asymmetry metrics, we identify six unambiguously settled host galaxies, demonstrating for the first time that a disturbed Hi morphology is not a universal feature of FRB host galaxies. However, we find another six that show clear signs of disturbance, one borderline case, and three which require deeper or more targeted observations to reach a conclusion; this brings the confirmed ratio of disturbed-to-settled FRB hosts to 11:6. Given that roughly a 1:1 ratio is expected for random background galaxies of similar type, our observed ratio yields a p-value of 0.222. Therefore, we conclude that contrary to earlier indications, there is no statistically significant excess of Hi disturbance in this sample of FRB host galaxies with respect to the general galaxy population, and hence we find no evidence for a fundamental connection between FRB progenitor formation and merger-induced star formation activity.
@@ 1 Introduction
Fast radio bursts (FRBs) are energetic pulses of radio emission that generally last for a few milliseconds in timescale Lorimer2007 (53). Whilst their extragalactic origins are well established Chatterjee2017 (14, 5), their generation mechanisms remain unclear, despite significant advances in the field. Dozens of progenitor models have been proposed to explain the phenomena, from outbursting neutron stars to interacting black holes (Platts2019, 71, 70, for reviews, see). The former, specifically young, highly magnetised neutron stars or magnetars, are favoured within the community due to observations of energetic FRBs found to be embedded within persistent radio sources (PRSs) (Metzger2017, 64, e.g.). This link was strengthened in 2020 with the detection of an FRB from a magnetar within our galaxy Andersen2020 (18, 10), although its isotropic-equivalent energy falls nearly 30 times lower than the next weakest FRB.
@@
Since their discovery, the population of detected FRBs has rapidly expanded, primarily due to dedicated observations from the Australian Square Kilometre Array Pathfinder (ASKAP, 42, 5, 80, ASKAP;), the Canadian Hydrogen Intensity Mapping Experiment (CHIME, 4, 17, CHIME;), MeerKAT MeerKAT (46, 7), and the Deep Synoptic Array (Kocz2019, 49, 73, 21, DSA;). Crucially, improvements to spatial resolution have allowed many FRBs to be localised to arcsecond or better accuracy, allowing for deep follow up photometry and spectroscopy of their host galaxies and environments (Bannister2019, 5, 40, 58, 8, 34, 77, 81, 16, e.g.). Such localisations are key to revealing the environments of FRBs and unlocking their potential for cosmological studies (McQuinn2014, 63, 55, 44, 21, e.g.).
@@
As of September 2025, over 100 host galaxies have been identified, spanning redshifts out to 2.15. These galaxies tend to be star-forming spiral galaxies, though most other galactic parameters such as stellar mass and mass-weighted age are highly varied Bhandari2022 (8, 34). These observations broadly support a magnetar progenitor hypothesis, as magnetars are expected to trace star formation through a core-collapse supernovae generation channel Metzger2017 (64). However, a number of counterexamples, such as a localisation to a globular cluster Kirsten2022 (48), and several to quiescent galaxies Eftekhari2025 (28, 79), muddy the waters. Furthermore, no statistically significant trend differing the hosts of repeating and apparently non-repeating FRBs has been observed Gordon2023 (34, 81), despite significant evidence suggesting fundamental differences in their respective origins ChimeC1 (17).
@@
Beyond analysis of the optical output of these host galaxies, a number of studies have been conducted on their atomic hydrogen (Hi) gas emission Michalowski21 (66, 47, 31, 52, 30). At redshifts greater than 0.1, the Hi line becomes weak and contaminated by RFI; with much of the localised FRB population lying beyond this redshift, only six host galaxies (excluding the Milky Way) have thus far been probed. All six of these galaxies have been detected in Hi, an expected result considering the star-forming nature of most hosts as derived from optical observations. Interestingly, a trend has emerged, with five of six of these detections exhibiting either a strongly asymmetric Hi spectrum or a highly disturbed spatial gas distribution. The single apparently symmetric profile, that of FRB 20211127I, was analysed with a low resolution and low signal-to-noise ratio (SNR) ASKAP observation, and thus its lack of disturbance was considered inconclusive Glowacki_wallaby (31).
@@
Such features, which are well documented in the literature (Holwerda2011, 41, 29, 11, 74, e.g.), are attributed to recent or ongoing disruptions to the galactic environment. The gaseous disk is sensitive to both external and internal perturbative mechanisms, including tidal interactions and mergers, ram pressure stripping, and feedback from supernovae or active galactic nuclei (Yu2022, 93, AGN;). Disturbed Hi distributions are not uncommon; many studies have found that a significant fraction of galaxies exhibit some sign of disturbance. Historical studies using single dish instruments focused on the asymmetry of the 1D profile; early estimates found that at least 50% of galaxies exhibit spectra with some level of asymmetry Richter94 (76, 38, 60). However, more recent studies use more stringent cutoffs; including these results Espada2011 (29, 88, 74, 93) and reevaluating the historical ones find anywhere between 10-40% of galaxies exhibit asymmetric Hi line profiles.
@@
In comparison, few studies have investigated similar properties in the spatial distribution of resolved Hi galaxies Holwerda2011 (41, 86). Reynolds2020LVHIS (75) recently probed 120 galaxies from the LVHIS LVHIS (50), VIVA VIVA (19), and HALOGAS HALOGAS (39) surveys, finding similar rates of asymmetry. However, with a number of more advanced interferometric observatories currently in the process of conducting large Hi surveys of resolved galaxies — such as the Widefield ASKAP L-band Legacy All-sky Blind surveY (WALLABY, 51, WALLABY,) on ASKAP, the Hi project within the MeerKAT International GigaHertz Tiered Extragalactic Exploration survey (MIGHTEEHI, 56, MIGHTEE-HI,), and the Apertif HI surveys ApertifHI (1) on the Westerbork Synthesis Radio Telescope (WRSTapertif, 85, WRST,) — many more observations will be made to properly pin down morphological asymmetry in Hi disks.
@@
Overall, the exact fraction of disturbed galaxies measured by each study is highly dependent on the samples used (e.g., isolated vs galaxy pairs), the observational properties (e.g. resolution and sensitivity), the method of quantification, and the point of cutoff in these methods above which a profile is considered disrupted. However, it can generally be said that disturbances are common, and are likely present in up to half of the galaxy population.
@@
Thus, given the low number of hosts probed, the early trend towards asymmetry in the population — with a ratio of 5:0 or 5:1 — is not necessarily significant at this stage. However, if this trend perseveres as more host galaxies are observed in Hi, it may indicate a connection between FRBs and a recent enhancement of star formation due to galactic interaction Michalowski21 (66). Indeed, studies over the past few decades have identified higher star formation rates in merging galaxies (Alonso2004, 3, 22, e.g.). Such a scenario would lend further credence to so called ‘fast track’ FRB progenitors such as magnetars, which are expected to trace star formation.
@@ 2.1 FRB Host Galaxy Targets
The FRBs selected for this study are taken from published datasets from the CRAFT survey (CRAFT, 54, 5, 80, Commensal Real-time ASKAP Fast Transients;), the CHIME/FRB Collaboration ChimeC1 (17, 16), and the MeerTRAP survey (MeerTRAP, 7, 72, 26, More TRAnsients and Pulsars;).
We place two criteria on targets for inclusion in this study. The first is that the host association is robust, which is evaluated through the PATH framework (Aggarwal2021, 2, Probabilistic Association of Transients to their Hosts;). PATH employs a Bayesian approach to calculate the probability a given transient event is associated with a galaxy, based on the event’s localisation and the galaxy’s magnitude and size. Following the convention used in Bhandari2022 (8, 34), we consider PATH posterior probabilities P(O—) 0.9 to be robust host associations.
Our second criterion is that the redshifts of the target host galaxies are . This is due to both the weak intensity of the Hi line, and also the prevalence of contaminating radio frequency interference (RFI) below 1300 MHz (0.092), largely due to the Global Navigation Satellite System (GNSS).
Based on these cuts, our sample consists of nine CHIME FRBs, six CRAFT FRBs, and two MeerTRAP FRBs. All but one CHIME target (FRB 20200223B) are apparently non-repeating bursts (“non-repeaters” herein), and one of the CRAFT targets (FRB 20211127I) is the FRB previously analysed in Glowacki_wallaby (31). The details of their host galaxies are presented in Table 1.
@@
| FRB 20181220A | CHIME | 2MFGC 17440 | 348.6982 | 48.3421 | 0.02746 | chimefrbs (9) |
| FRB 20181223C | CHIME | SDSS J120340.98+273251.4 | 180.9207 | 27.5476 | 0.03024 | chimefrbs (9) |
| FRB 20190418A | CHIME | SDSS J042314.96+160425.6 | 65.8123 | 16.0738 | 0.07132 | chimefrbs (9) |
| FRB 20190425A | CHIME | UGC 10667 | 255.6625 | 21.5767 | 0.03122 | chimefrbs (9) |
| FRB 20200223B† | CHIME | SDSS J003304.68+284952.6 | 8.2695 | 28.8313 | 0.06024 | Ibik2023 (43) |
| FRB 20200723B | CHIME | NGC 4602 | 190.1538 | -5.1328 | 0.0085 | frb20200723b (82) |
| FRB 20201123A | MeerTRAP | J173438.35-504550.4 | 263.6596 | -50.7641 | 0.0507 | Rajwade2022 (72) |
| FRB 20210405I | MeerTRAP | 2MASS J1701249-4932475 | 255.3537 | -49.5465 | 0.0656 | frb20210405 (27) |
| FRB 20211127I | CRAFT | WALLABY J131913–185018 | 199.8082 | -18.8378 | 0.04695 | Glowacki_wallaby (31) |
| FRB 20211212A | CRAFT | SDSS J102924.22+012139.2 | 157.3509 | 1.3608 | 0.0715 | Shannon2024 (80) |
| FRB 20231229A | CHIME | UGC 1234 | 26.4658 | 35.1110 | 0.0190 | Amiri2025 (16) |
| FRB 20231230A | CHIME | J045109.39+022205.02 | 72.7887 | 2.3681 | 0.0298 | Amiri2025 (16) |
| FRB 20240201A | CRAFT | WISEA J095937.44+140519.4 | 149.9060 | 14.0887 | 0.04273 | Shannon2024 (80) |
| FRB 20240210A | CRAFT | WISEA J003506.47–281619.1 | 8.7768 | -28.2723 | 0.02369 | Shannon2024 (80) |
| FRB 20240312D | CRAFT | WISEA J032629.50-543557.2 | 51.6227 | -54.5993 | 0.04973 | Gordon2025 (33) |
| FRB 20240615B | CRAFT | WISEA J021551.26-143729.8 | 33.9633 | -14.6247 | 0.0730 | Gordon2025 (33) |
| FRB 20250316A | CHIME | NGC 4141 | 182.4472 | 58.8492 | 0.00635 | Amiri2025 (16) |
@@ 2.2 Hi Observations & Processing
We first searched for archival Hi datasets, and found observations of the hosts of FRB 20231229A and FRB 20250316A Amiri2025 (16). Both galaxies have been observed in Hi twice; we take the most recent spectra available in each case, corresponding to an ALFALFA (Haynes2018, 37, Arecibo Legacy Fast ALFA;) survey observation and an Effelsberg 100-m observation Haynes1999 (36) respectively. Both of these are at spectral resolutions of 11 km s-1, with single-dish beams encompassing several times the angular size of the galaxy.
A search of the MeerKAT archive hosted by the South African Radio Astronomical Observatory (SARAO) revealed that the coordinates of three targets were contained in archival band observations. The host of FRB 20200723B frb20200723b (82) was the target of a 0.5 hour 4K channelisation (corresponding to 44 km s-1resolution) pointing from ”Part 2: 1.28 GHz MeerKAT survey of 12MGS Galaxies in the Southern Hemisphere”, project ID SCI-20220822-LM-01, capture block ID 1671245181. The host of FRB 20201123A Rajwade2022 (72) was contained in a 1 hour 4K channelisation pointing by the ”Legacy Survey of the Galactic Plane”, project ID SSV-20180505-FC-01, capture block ID 1610418665, a distance of 0.4 from the phase centre. The host of FRB 20231230A Amiri2025 (16) was contained in a 7.5 hour 4K channelisation pointing from the proposal ”Observation of the galaxy cluster A520”, project ID SCI-20210212-SG-01, capture block ID 1638727749, a distance of 0.9 from the phase centre.
The remaining targets had no other archival data available, and thus we pursued targeted observation with the upgraded Giant Metrewave Radio Telescope (GMRT, 84, 35, uGMRT;) and the Five-hundred-meter Aperture Spherical Telescope (FAST, 68, FAST;) for northern hemisphere targets, and MeerKAT for southern hemisphere targets. We present summaries of all Hi observations in Table 2.
@@
| FRB 20240210A | MeerKAT | 28th July 2025 | 26.123 | 240 | J0408-6545 | J0025-2602 |
| FRB 20240312D | MeerKAT | 29th June 2025 | 26.123 | 240 | J0408-6545 | J0303-6211 |
| FRB 20240615B | MeerKAT | 3rd August 2025 | 26.123 | 240 | J0408-6545 | J0240-2309 |
| FRB 20250316A | Effelsbergd | Oct/Nov 1989 | 11.965 | - | - | - |
@@
Haynes2018 (37) - ALFALFA survey was conducted over a 7 year time span with an average integration time of 48 seconds.
@@
Haynes1999 (36) - precise on source time unknown.